What is the force applied when mass of the body is 5kg and acceleration is 1000cm per second square ?

Grade:12th pass

3 Answers

DIWAKAR PANDEY
6 years ago
Dear its simple Just convert cm/s2 into m/s2 So it become a=10m/s2 So f=ma F=5×10= 50N Hope it is now clear......
shubhankit
6 years ago
1000cm/second ^2=10m/second^2Hence,force=mass×acc.Also mass =5kg according to question, I.e.=5×10=50NHope it helps"
naveen
6 years ago
Hello friends in cgs system mass in gm and acceleration in cm per sec square soConvert mass into gm 5000 into 1000 equals to 5000000 dynes (Or)In si units kg in mass and m per sec square so f=m×g 5×10=50newtons